A neutron causes 90232Th to change according to the following reaction: 01n + 90232Th ? ZAX + ? The ZAX nucleus subsequently undergoes ?- decay, and its daughter does too. What is the atomic mass number of the final nucleus?
Answer #1


the given nuclear reaction is,


(Z=0)(A=1)n+(Z=90)(A=232)Th---->(Z=?)(A=?)X+(Z=0)(A=0)gamaa


equate the sum of atomic mass number(A) both sides

1+232=A+0

A=233

and

equate the sum of atomic number(Z) both sides

0+90=Z+0

Z=90


hence

Atomic mass number of the final nucleus is A=233

Atomic number of the final nucleus is Z=90